OpenStudy (anonymous):

A 6V battery is connected to the two plates of a capacitor which is separated by o.30 cm. Find the magnitude of the electric field between the plates. Derive any equation used. Could someone plz help with derivation and kindly explain it. thx

OpenStudy (anonymous):

For a constant force, work is F*d = qE*d since F = qE. Now work is also qV... so setting the two expressions for work equal we find V = Ed, or E = V/d (units are volts/m) Plug in and solve, making sure to convert cm to m